(14 pts] a) Determine the critical value ze for a level of confidence of 60%. In other words, find z, such that P(-z, <z < z) please show work
Answer #1

Answer: 12 _ a) Guiven level of confidence oo% PC-Zc LZ< Zc) = 0.60 k_o > Il/a zidla Here, x=1-0.60 = 0.40 -0.40 = 0:20 2 -|-= 0,4 = 0.20 1 = 1-04 = 0.80 Here the we have to see the probability 0.20 ond_0.80 in normal table ond wolde down I-score co